In a bolt manufacturing factory, machines A, B, and C produce 25%, 30% and 45%
of the total output, respectively. Of their outputs, 7%, 6% and 4% are defective bolts,
respectively. If a bolt drawn at random from the production is found to be defective,
what is the probability that it was manufactured by machine C?
1
Expert's answer
2020-06-03T18:39:40-0400

A : Bolt Manufactured by machine A

B : Bolt Manufactured by machine B

C : Bolt Manufactured by machine C

D : Bolt is defective

We need to find the probability that the bolt is manufactured by machine C , if it is defective i.e. "P(C\/D)"

"P(A) = 0.25," "P(D\/A)=0.07,"



"P(B) = 0.30," "P(D\/B)=0.06"



"P(C) = 0.45," "P(D\/C)=0.04"



"P(C / D) = \frac{P( D / C) P(C) }{P(A) P(D / A)+P(B)P(D / B)+P(C)P(D /C)}"



"P(C/D) = \frac {(0.04).(0.45)}{(0.25).(0.07)+(0.30).(0.06)+(0.45).(0.04)} = \frac{36}{107}"